OpenStudy (anonymous):

Solve this system of equations. Write the answer as an ordered pair. -5x - 2y = -2 x - 7y = -7

OpenStudy (anonymous):

i barely know substitution

OpenStudy (haseeb96):

it is too much easy equations just separate the x from equation 2 and put it in 2 and get y same as it is again resulting y put in equation 2 and get x

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

dumb it down for me

OpenStudy (haseeb96):

x=-7 + 7 y put it in equation 1 -5(-7 + 7y) -2y =-2 35 -35y -2y =-2 -37y=-37 y=1 now put it in equation 1 and get x

OpenStudy (anonymous):

oh, ok thank you

OpenStudy (anonymous):

(1,-7)?